Solve the system of equations: 2x+y=5
-3x+2y=-1

Answer by mananth(12270)

2x+y=5 ................1
-3x+2y=-1..............2
..
multiply equation 1 by -2 and add it to equation 2
-4x-2y-3x+2y=-10-1
-7x=-11
x= 11/7
..
plug value of x in equation 1
2*11/7 +y= 5
y= 5- 22/7
y=13/7